# Generalized local jocabians and commutative group stacks

### v2

@@ 3. Grothendieck’s generalized local jacobians
In this section we provide $\mathbb{H}^{alg}(X)$ with a natural filtration, induced by the standard filtration by codimension on cohomology. The associated graded to this filtration possesses a description in local terms involving local cohomology. When working with smooth schemes over $k$, the local factors will directly related to Grothendieck’s orginial construction of local generalized jacobians from his letter to Serre (letter from 9.8.1960, [CS01] page 109).
@@ 3.1. The filtration by (co)dimension
We let $d\geq 0$ be an integer, and consider $Op_{d}(X)$ the poset of opens $U\subset X$ containing $X^{(\geq d)}$. An open $U$ belongs to $Op_{d}(X)$ if and only if the closed complement $U-Z$ has dimension at most $d-1$. The only element in $Op_{0}(X)$ is $X$ itself, whereas $\emptyset\in Op_{dimX+1}(X)$, and obviously $Op_{d}(X)\subset Op_{d+1}(X)$. We define $F_{d}\mathbb{H}^{alg}(X)$ to be
@@
Proof. For any inclusion of open $V\subset U$, with $U\in Op_{d}(X)$ and $V\in Op_{d+1}(X)$, the closed complement $U-V$ only has a finite number of points of dimension $d$. Indeed, if $(U-V)^{(d)}$ is infinite, then $U-V$ must be of dimension at least $d+1$ and thus must contain at least one point of dimension $d+1$, which contradicts $V\in Op_{d+1}(X)$. Note also that for $x\in(U-V)^{(d)}$, we have $X_{x}\cap V=X_{x}-\{x\}$, as $x$ is the only point of $X_{x}$ of dimension $d$ in $X$. We thus get this way a commutative square of schemes
[…]
and thus an induced commutative square on the vertical cofibers on algebraic homologies
@@
are costacks with values in $Pro(\mathbf{AbSt}_{k})$ on the small Zariski site $X_{Zar}$. It is therefore enough to check that the above morphism induces an equivalence on the fibers at a given point $x\in X$. This fiber is clearly 00 if $x$ is not of codimension $d$. If $x$ is of codimension $d$ the fiber of the right hand side is equivalent to
@@
$co\lim_{x\in U}\mathbf{Map}(U,G)\to\mathbf{Map}(X_{x},G)\qquad co\lim_{x\in U}\mathbf{Map}(U-Z_{x},G)\to\mathbf{Map}(X_{x}-\{x\},G)$
@@
As we already saw, the commutative algebraic group stacks are not stable by taking cofibers. So, a priory $Gr_{d}\mathbb{H}^{alg}(X)$ merely belongs to $Pro(\mathbf{AbSt}_{k}^{lalg})$, the $\infty$-category of pro-objects in commutative locally algebraic groups stacks. However, it is easy to see by universal property that for any connected smooth variety $U$ over $k$, the group of connected components of $\mathbb{H}_{0}^{alg}(U)$ is canonically isomorphic to $\mathbb{Z}$, as any morphism from $U$ to a discrete group $G$ must be constant. As a result, for any morphism $U\to V$ of such varieties, the induced morphism $\mathbb{H}_{0}^{alg}(U)\to\mathbb{H}_{0}^{alg}(V)$ is an isomorphism on the group of connected components. This easily implies that $Gr_{d}\mathbb{H}^{alg}(X)$ is a pro-object in $\mathbf{AbSt}_{k}$.
@@
Let $X$ be a connected and smooth $k$-scheme of dimension $n$. Then, for all integer $d$ the object $Gr_{d}\mathbb{H}^{alg}(X)$ is $(n-d)$-connective: its pro-homotopy groups $\pi_{i}(Gr_{d}\mathbb{H}^{alg}(X))$ vanish for $i<n-d$.
Proof. When $d=n$ there is nothing to prove as by construction all $Gr_{d}\mathbb{H}^{alg}(X)$ are 00-connective. For $d<n$ we have to show that $\pi_{i}(Gr_{d}\mathbb{H}^{alg}(X))=0$ for $i<n-d$. Equivalently we have to prove that any $G\in\mathbf{AbSt}_{k}^{alg}$ which is $(n-d-1)$-truncated, then $\mathbf{Map}_{\mathbf{AbSt}_{k}}(Gr_{d}\mathbb{H}^{alg}(X),G)\simeq*$. By proposition 3.2 this is equivalent to say that $\mathbb{H}_{x}(X,G)\simeq 0$ for all point $x\in X$ of dimension $d$. We then proceed by Postnikov decomposition on $G$, and we thus reduce the corollary to the following two statements.
@@
    For all $x\in X^{(d)}$, and all abelian group scheme $G$, separated and locally of finite type, we have $H_{x}^{0}(X,G)=0$.
@@
    For all $x\in X^{(d)}$ and all abelian group scheme $G$, separated and of finite type, we have $H^{i}_{x}(X,G)=0$ for all $i<n-d$
The first of this statement is obviously true because $d<n$ and thus $X_{x}-\{x\}$ is dense in $X_{x}$. The second statement has already been reminded in proposition 1.2. $\Box$
@@ 3.2. Grothendieck’s construction
We will now compare the general notion of algebraic homology with the construction sketched by Grothendieck in [CS01, Page 109] (which also appears in a letter to L. Breen, see [Gro, cote 134-2] page 60). As we are working with the fppf topology, we will have to modify slightly the presentation that was originally done by Grothendieck in the context of the Zariski topology. For this, and all along this section, we will assume that $X$ is a smooth scheme, connected and of dimension $n$ over $k$. Let $x\in X^{(d)}$ be a point of dimension $d$ on $X$.
We consider the functor $G\mapsto H_{x}^{n-d}(X_{x},G)$, of (fppf) local cohomology at $x$, defined on the category of abelian group schemes (separated and of finite type) over $k$. By purity (proposition 1.2) we have $H_{x}^{n-d-1}(X_{x},G)\simeq 0$ and thus the above functor is left exact in $G$. It is therefore pro-representable by an abelian pro-gourp scheme $J_{x}(X)$.
For an integer $d$ we then set
@@
which is a pro object in the category of commutative group schemes locally of finite type over $k$. We denote by $J_{*}(X)$ the graded object $\oplus_{m}J_{m}(X)$. It comes equipped with a homological differential $J_{m}(X)\to J_{m-1}(X)$ defined as follows. For a fixed $z\in X^{(n-m+1)}$, we define a morphism
@@
by defining a natural transformation on the functor they pro-represent. The right hand side pro-represents the functor $G\mapsto H^{m-1}_{z}(X_{z},G)$, whereas the left hand side pro-represents the functor $G\mapsto\oplus_{x\in X^{(n-m)}}H^{m}_{x}(X_{x},G)$. In order to define $d_{z}$ we thus have to construct a morphism on local cohomologies
@@
functorial in $G$. This morphism is the usual differential in the Cousin complex (see [CTHK97, §1]). Any element $\alpha\in H^{m-1}_{z}(X_{z},G)$ comes from an element $\alpha_{U}\in H^{m-1}_{Z}(U,G)$ for an open $U\subset X$ containing $z$, and where $Z$ is the closure of $z$ in $U$. We denote by $x_{1},\dots,x_{q}$ the points of $X$ of dimension $n-m$ which are specializations of $z$ but not belonging to $U$, and we let $T=X-U$. We consider the boundary map $H^{m-1}(U,G)\to H^{m-1}_{T}(X,G)$, together with the restriction morphism $H^{m}_{T}(X,G)\to H^{m}_{x}(X_{x},G)$ along the inclusion of pairs $(X_{x},\{x\})\subset(X,T)$. We get this way a composed morphism
@@
    The pro-algebraic group $J_{d}(X)=\prod_{x\in X^{(n-i)}}J_{x}(X)$ is called Grothendieck’s local generalized jacobian of $X$ at $x$.
@@
    The complex of local jacobians, with the differentials define above, will be denoted by
@@
As a first relation between the complex $J_{*}(X)$ and algebraic homology, we have the following direct observation, which is a direct consequence of corollary 3.4, proposition LABEL:p5 and and the definitions of $J_{*}(X)$.
@@
$\pi_{n-d}(Gr_{n-d}\mathbb{H}^{alg}(X))\simeq J_{d}(X).$
The above result can be made more precise. Indeed, the filtration by dimension on $\mathbb{H}^{alg}(X)$ induces a spectral sequence
$\pi_{p}(Gr_{q}\mathbb{H}^{alg}(X))\Rightarrow\pi_{p+q}(\mathbb{H}^{alg}(X)),$
and under the identification of the corollary the differential in the complex $J_{*}(X)$ corresponds to the differential on the $E_{1}$-term of the spectral sequence. Therefore, by the purity 3.4, Grothendieck’s local generalized jacobian $J_{d}(X)$ appear as the lowest non-trivial homotopy group of $Gr_{(n-d)}\mathbb{H}^{alg}(X)$. However, the graded pieces $Gr_{(n-d)}\mathbb{H}^{alg}(X)$ are in general not concentrated in a single degree, and thus $Gr_{(n-d)}\mathbb{H}^{alg}(X)$ and $J_{d}(X)$ differs. We will now see however that the complex $J_{*}(X)$ is essentially enough to recover the cohomology of $X$ with coefficients in certain commutative algebraic group as stated in [CS01, Page 109].
The corollary 3.6 can also be phrased by stating that the complex of pro-algebraic groups $J_{*}(X)$ is the 00-truncation of the filtered object $\mathbb{H}^{alg}(X)$ with respect to the Beilinson’s t-structure on filtered objects. As a result, we do have a canonical projection of filtered objects
@@
where the filtration on the right hand side is the naive filtration $J_{*\leq n-d}(X)\subset J_{*}(X)$. This projection is not an equivalence, as the graded pieces $Gr_{d}\mathbb{H}^{alg}(X)$ are not purely concentrated in a single degree in general. However, the projection $q$ induces an isomorphisms on certain ext groups showing that the complex $J_{*}(X)$ can be used in order to compute the individual cohomology groups $H^{i}_{fppf}(X,G)$ with coefficient in a commutative algebraic group $G$ (affine if $i>0$, and unipotent if $i>1$). In order to explain this, we need to first introduce a comparison functor, between the naive derived category of commutative group schemes and our category of commutative derived group stacks.
We know that the category $\mathbf{AbSch}_{k}$, of separated commutative group schemes of finite type over $k$ is an abelian category. Its category of pro-objects $Pro(\mathbf{AbSch}_{k})$ is known to be abelian and to possess enough projective objects (see [Mil70]). As such it possesses a bounded derived category $\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))$, obtained by localizing the category of bounded complexes in $Pro(\mathbf{AbSch}_{k})$ with respect to the quasi-isomorphisms. Because the functor sending $G\in Pro(\mathbf{AbSch}_{k})$ to the fppf-sheaf it represents is exact, there is a canonical $\infty$-functor
@@
It is well known that this functor is not fully faithful in general, as this can be seen on higher ext-groups. The ext-groups defined in $\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))$ are the naive extension groups of commutative pro-algebraic groups, and can be described in terms of Yoneda extensions. On the other side, the extensions groups in $\mathbf{AbSt}_{k}$ are the ext-groups of sheaves. It is know that, when $k$ is algebraically closed, $Ext^{i}$ vanish when $i>2$ and when computed in $\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))$ (see [Mil70]), but this is not the case in $\mathbf{AbSt}_{k}$ (see for instance [Bre69]).
@@
$\phi^{i}_{G}:[J_{*}(X),G[i]]_{\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))}\longrightarrow[h(J_{*}(X)),h(G)[i]]\longrightarrow[\mathbb{H}^{alg}(X),h(G)[i]].$
@@
Proof. As in [Mil70] we can use Galois descent in order to reduce to the case where $k$ is algebraically closed. We start by using that the projection $q:\mathbb{H}^{alg}(X)\to J_{*}(X)$ is compatible with the filtrations on both sides: the filtration by dimension on $\mathbb{H}^{alg}(X)$ and the naive filtration on $J_{*}(X)$. These filtrations induce convergent spectral sequences
$[Gr_{n-q}\mathbb{H}^{alg}(X),G[p]]\Rightarrow H^{p+q}_{fppf}(X,G)\qquad Ext^{p}_{naive}(J_{q}(X),G)\Rightarrow Ext^{p+q}_{naive}(J_{*}(X),G).$
The morphism $\phi_{G}$ of the proposition is then compatible with these spectral sequences, and we are thus reduced to show that, for fixed $p$ and $q$, the induced morphism
$Ext^{p}_{naive}(J_{q}(X),G)\to[Gr_{n-q}\mathbb{H}^{alg}(X),G[p]]$
is an isomorphism. By proposition 3.2, the right hand side can be identified with local cohomology $\coprod_{x\in X^{(n-q)}}H^{p}_{x}(X,G)$. We also note that by construction $J_{q}(X)=\prod_{x\in X^{(n-q)}}J_{x}(X)$, and thus we are reduced to show that for a given point $x\in X$, of codimension $d$, and all $i\geq 0$, the natural morphism
@@
under the extra condition stated in the proposition: $G$ is affine if $d+i>0$ and unipotent if $d+i>1$. Recall here that $J_{x}(X)$ corepresents the functor $G\mapsto H^{d}_{x}(X,G)$, where $d$ is the codimension of $x$ in $X$.
We first deal with the two special cases $i+d=0$ and $i+d=1$. When $i+d=0$, and thus $i=d=0$, we have $Ext^{0}(J_{x}(X),G)=Hom(J_{x}(X),G)\simeq H^{d}_{x}(X,G)$ and it is then obvious that $\psi^{0}_{G}$ is bijective. Suppose now that $d+i=1$, and thus $G$ is assumed to be affine. If $x$ is of codimension $1$, we have $i=0$, and again $\psi_{G}^{0}$ is bijective by definition of $J_{x}(X)$. If $x$ is of codimension 00 and $i=1$, we have to prove that the natural morphism
@@
where $K(X)$ is the fraction field of $X$. To show this we start by the following lemma treating the case of $G=\mathbb{G}_{a}$ and $G=\mathbb{G}_{m}$.
@@
With the notation above we have
$\psi_{G}^{1}:Ext^{1}_{naive}(J_{0}(X),\mathbb{G}_{a})\simeq Ext^{1}_{naive}(J_{0}(X),\mathbb{G}_{m})\simeq 0.$
[…]
@@
$0\to G\to\mathbb{G}_{m}\to\mathbb{G}_{m}\to 0\qquad 0\to G\to\mathbb{G}_{a}\to\mathbb{G}_{a}\to 0$
We then consider the action of $\psi_{G}^{1}$ on the associated long exact sequences, and lemma 3.8 shows that indeed $\psi_{G}^{1}$ is bijective for any $G$.
We now turn to the case where $i+d>1$ and $G$ is unipotent, which is proven in a very similar fashion than the case $i=1$ and $d=0$. To begin with, we prove that $\psi_{G}^{i}$ is an isomorphism for $G=\mathbb{G}_{a}$. Indeed, when $i=0$ this is obvious by the definition of $J_{x}(X)$. For $i>0$ the domain and codomain of $\psi_{\mathbb{G}_{a}}^{i}$ vanishes. Indeed, $H^{d+i}_{x}(X,\mathbb{G}_{a})\simeq H^{d+i}_{x}(X,\mathcal{O})\simeq 0$ for $i>0$ by local duality for coherent sheaves. On the other hand, $Ext^{i}_{naive}(J_{x}(X),\mathbb{G}_{a})=0$ by the following lemma.

### v3

@@ 3. Grothendieck’s generalized local jacobians
In this section we provide $\mathbb{H}^{alg}(X)$ with a natural filtration, induced by the standard filtration by codimension on cohomology. The associated graded to this filtration possesses a description in local terms involving local cohomology. When working with smooth schemes over $k$, the local factors will be related to Grothendieck’s original construction of local generalized jacobians from his letter to Serre (letter from 9.8.1960, [10] page 109).
@@ 3.1. The filtration by (co)dimension
We let $d\geq 0$ be an integer, and consider $Op_{d}(X)$ the poset of opens $U\subset X$ containing $X^{(\geq d)}$. An open $U$ belongs to $Op_{d}(X)$ if and only if the closed complement XUX-U has dimension at most d1d-1. The only element in $Op_{0}(X)$ is $X$ itself, whereas $\emptyset\in Op_{dimX+1}(X)$, and obviously $Op_{d}(X)\subset Op_{d+1}(X)$. We define $F_{d}\mathbb{H}^{alg}(X)$ to be
@@
Proof. For any inclusion of open $V\subset U$, with $U\in Op_{d}(X)$ and $V\in Op_{d+1}(X)$, the closed complement UVU-V only has a finite number of points of dimension $d$. Indeed, if $(U-V)^{(d)}$ is infinite, then UVU-V must be of dimension at least $d+1$ and thus must contain at least one point of dimension $d+1$, which contradicts $V\in Op_{d+1}(X)$. Note also that for $x\in(U-V)^{(d)}$, we have $X_{x}\cap V=X_{x}-\{x\}$, as $x$ is the only point of $X_{x}$ of dimension $d$ in $X$. We thus get this way a commutative square of schemes
${\coprod_{x\in(U-V)^{(d)}}X_{x}-\{x\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\coprod_{x\in(U-V)^{(d)}}X_{x}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U,}$
and thus an induced commutative square on algebraic homologies, which induces a map on the corresponding vertical cofibers
@@
are costacks with values in $Pro(\mathbf{AbSt}_{k})$ on the small Zariski site $X_{Zar}$. It is therefore enough to check that the above morphism induces an equivalence on the fibers at a given point $x\in X$. This fiber is clearly 0 if $x$ is not of dimension $d$. If $x$ is of dimension $d$ the fiber of the right hand side is equivalent to
@@
$\underset{x\in U}{colim}\mathbf{Map}(U,G)\to\mathbf{Map}(X_{x},G)\qquad\underset{x\in U}{colim}\mathbf{Map}(U-Z_{x},G)\to\mathbf{Map}(X_{x}-\{x\},G)$
@@
As we already saw, the commutative algebraic group stacks are not stable by taking cofibers. So, a priory $Gr_{d}\mathbb{H}^{alg}(X)$ merely belongs to $Pro(\mathbf{AbSt}_{k}^{lalg})$, the $\infty$-category of pro-objects in commutative locally algebraic groups stacks. However, it is easy to see by universal property that for any connected smooth variety $U$ over $k$, the group of connected components of $\mathbb{H}_{0}^{alg}(U)$ is canonically isomorphic to $\mathbb{Z}$, as any morphism from $U$ to a discrete group $G$ must be constant. As a result, for any morphism $U\to V$ of such varieties, the induced morphism $\mathbb{H}_{0}^{alg}(U)\to\mathbb{H}_{0}^{alg}(V)$ is an isomorphism on the group of connected components. This easily implies that $Gr_{d}\mathbb{H}^{alg}(X)$ is a pro-object in $\mathbf{AbSt}^{alg}_{k}$.
@@
Let $X$ be a connected and smooth $k$-scheme of dimension $n$. Then, for all integer $d$ the object $Gr_{d}\mathbb{H}^{alg}(X)$ is (nd)(n-d)-connective: its pro-homotopy groups $\pi_{i}(Gr_{d}\mathbb{H}^{alg}(X))$ vanish for i<ndi<n-d.
Proof. When $d=n$ there is nothing to prove as by construction all $Gr_{d}\mathbb{H}^{alg}(X)$ are 0-connective. For $d<n$ we have to show that $\pi_{i}(Gr_{d}\mathbb{H}^{alg}(X))=0$ for i<ndi<n-d. Equivalently we have to prove that any $G\in\mathbf{AbSt}_{k}^{alg}$ which is (nd1)(n-d-1)-truncated, then $\mathbf{Map}_{\mathbf{AbSt}_{k}}(Gr_{d}\mathbb{H}^{alg}(X),G)\simeq*$. By the proposition 3.2 this is equivalent to say that $\mathbb{H}_{x}(X,G)\simeq 0$ for any point $x\in X$ of dimension $d$. We then proceed by Postnikov decomposition on $G$, and we thus reduce the corollary to the following two statements.
@@
    For all $x\in X^{(d)}$, and all abelian group scheme $G$ locally of finite type, we have $H_{x}^{0}(X,G)=0$.
@@
    For all $x\in X^{(d)}$ and all abelian group scheme $G$ of finite type, we have $H^{i}_{x}(X,G)=0$ for all i<ndi<n-d
The first of this statement is obviously true because $d<n$ and thus $X_{x}-\{x\}$ is dense in $X_{x}$. The second statement has already been reminded in the proposition 1.2. $\Box$
@@ 3.2. Grothendieck’s construction
We will now compare the general notion of algebraic homology with the construction sketched by Grothendieck in [10, page 109] (which also appears in a letter to L. Breen, see [14, cote 134-2] page 60). As we are working with the fppf topology, we will have to modify slightly the presentation that was originally done by Grothendieck in the context of the Zariski topology. For this, and all along this section, we will assume that $X$ is a smooth scheme, connected and of dimension $n$ over $k$. Let $x\in X^{(d)}$ be a point of dimension $d$ on $X$.
We consider the functor $G\mapsto H_{x}^{n-d}(X_{x},G)$, (fppf) local cohomology at $x$, defined on the category of abelian group schemes (separated and of finite type) over $k$. By purity (Proposition 1.2) we have $H_{x}^{n-d-1}(X_{x},G)\simeq 0$ and thus the above functor is left exact in $G$. It is therefore pro-representable by an abelian pro-group scheme $J_{x}(X)$.
For an integer $m$ we define
@@
which is a pro-object in the category of commutative group schemes locally of finite type over $k$. We denote by $J_{*}(X)$ the graded object $\oplus_{m}J_{m}(X)$. It comes equipped with a homological differential $J_{m}(X)\to J_{m-1}(X)$ defined as follows. For a fixed $z\in X^{(n-m+1)}$, we define a morphism
@@
by defining a natural transformation on the functor they pro-represent. The right hand side pro-represents the functor $G\mapsto H^{m-1}_{z}(X_{z},G)$, whereas the left hand side pro-represents the functor $G\mapsto\oplus_{x\in X^{(n-m)}}H^{m}_{x}(X_{x},G)$ (note that because $G$ is a pro-constant object we have $Hom(\prod_{x\in X^{(n-m)}}J_{x}(X),G)\simeq\oplus_{x\in X^{(n-m)}}Hom(J_{x}(X),G)$).
In order to define $d_{z}$ we thus have to construct a morphism on local cohomologies
@@
functorial in $G$. This morphism is the usual differential in the Cousin complex (see [9, §1]). Any element $\alpha\in H^{m-1}_{z}(X_{z},G)$ comes from an element $\alpha_{U}\in H^{m-1}_{Z}(U,G)$ for an open $U\subset X$ containing $z$, and where $Z$ is the closure of $z$ in $U$. We denote by $x_{1},\dots,x_{q}$ the points of $X$ of dimension nmn-m which are specializations of $z$ but not belonging to $U$, and we let T=XUT=X-U. We consider the boundary map $H^{m-1}(U,G)\to H^{m-1}_{T}(X,G)$, together with the restriction morphism $H^{m}_{T}(X,G)\to H^{m-1}_{x}(X_{x},G)$ along the inclusion of pairs $(X_{x},\{x\})\subset(X,T)$. We get this way a composed morphism
@@
    For any $x\in X$ of dimension nmn-m, the pro-algebraic group $J_{x}(X)$ is called Grothendieck’s local generalized jacobian of $X$ at $x$.
@@
    The pro-algebraic group $J_{m}(X):=\prod_{x\in X^{(n-m)}}J_{x}(X)$ called Grothendieck’s local generalized jacobian of $X$.
  3. (3)
    The complex of local generalized jacobians, with the differentials defined above, will be denoted by
@@
As a first relation between the complex $J_{*}(X)$ and algebraic homology, we have the following direct observation, which is a direct consequence of the proposition 3.4 and and the definitions of $J_{*}(X)$.
@@
$\pi_{d}(Gr_{d}\mathbb{H}^{alg}(X))\simeq J_{n-d}(X).$
By the purity result 3.4, Grothendieck’s local generalized jacobians $J_{d}(X)$ therefore appear as the lowest non-trivial homotopy group of $Gr_{(n-d)}\mathbb{H}^{alg}(X)$. However, the graded pieces $Gr_{(n-d)}\mathbb{H}^{alg}(X)$ are in general not concentrated in a single degree, and thus $Gr_{(n-d)}\mathbb{H}^{alg}(X)$ and $J_{d}(X)$ differ. We will now see however that the complex $J_{*}(X)$ is essentially enough to recover the cohomology of $X$ with coefficients in certain commutative algebraic groups as stated in [10, page 109].
The corollary 3.6 can also be phrased by stating that the complex of pro-algebraic groups $J_{*}(X)$ is the $n$-th truncation of the $n$-connective filtered object $\mathbb{H}^{alg}(X)$ with respect to the Beilinson’s t-structure on filtered objects. We remind that in our context of decreasing filtrations, this t-structure is characterized by the fact that the connective objects are the filtered objects $F_{*}E$ such that $Gr_{p}(E)$ is p-p-connective for all $p$: $\pi_{i}(Gr_{p}(E))=0$ for all i<pi<-p (see for instance [27]). As a result, we do have a canonical projection of filtered objects
@@
where the filtration on the right hand side is the naive filtration $F_{d}J_{*}(X):=J_{\geq n-d}(X)\subset J_{*}(X)$. This projection is not an equivalence, as the graded pieces $Gr_{d}\mathbb{H}^{alg}(X)$ are not purely concentrated in a single degree in general. However, the projection $q$ induces an isomorphism on certain ext groups showing that the complex $J_{*}(X)$ can be used in order to compute the individual cohomology groups $H^{i}_{fppf}(X,G)$ with coefficients in a commutative algebraic group $G$ (affine if $i>0$, and unipotent if $i>1$). In order to explain this, we need to first introduce a comparison functor, between the naive derived category of commutative group schemes and our category of commutative derived group stacks.
We know that the category $\mathbf{AbSch}_{k}$, of commutative group schemes of finite type over $k$ is an abelian category. Its category of pro-objects $Pro(\mathbf{AbSch}_{k})$ is known to be abelian and to possess enough projective objects (see [19]). As such it possesses a bounded derived category $\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))$, obtained by localizing the category of bounded complexes in $Pro(\mathbf{AbSch}_{k})$ with respect to the quasi-isomorphisms. Because the functor sending $G\in Pro(\mathbf{AbSch}_{k})$ to the fppf-sheaf it represents is exact, there is a canonical $\infty$-functor
@@
It is well known that this functor is not fully faithful in general, as this can be seen on higher ext-groups. The ext-groups defined in $\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))$ are the naive extension groups of commutative pro-algebraic groups, and can be described in terms of Yoneda extensions. On the other side, the extension groups in $\mathbf{AbSt}_{k}$ are the ext-groups of sheaves. It is know that, when $k$ is algebraically closed, $Ext^{i}$ vanish when $i>2$ and when computed in $\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))$ (see [19]), but this is not the case in $\mathbf{AbSt}_{k}$ (see for instance [4]).
@@
$\phi^{i}_{G}:[J_{*}(X),G[i]]_{\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))}\longrightarrow[h(J_{*}(X)),h(G)[i]]\longrightarrow[\mathbb{H}^{alg}(X),h(G)[i]],$
where we have denoted by [,][-,-] the morphism in the homotopy category of stacks
$F_{1},F_{2}]=\pi_{0}(Map_{\mathbf{St}_{k}}(F_{1},F_{2})).$
@@
Proof. As in [19] we can use Galois descent in order to reduce to the case where $k$ is algebraically closed. We start by using that the projection $q:\mathbb{H}^{alg}(X)\to J_{*}(X)$ is compatible with the filtrations on both sides: the filtration by dimension on $\mathbb{H}^{alg}(X)$ and the naive filtration on $J_{*}(X)$. To prove the proposition we will proceed by descending induction on $d$ and will prove that the induced morphism
$F_{d}(q):F_{d}\mathbb{H}^{alg}(X)\to J_{\geq n-d}(X)$
induces isomorphisms
$J_{\geq n-d}(X),G[i]]_{\mathcal{D}^{b}(Pro(\mathbf{AbSch}_{k}))}\simeq[F_{d}\mathbb{H}^{alg}(X),G[i]]$
for all $d$, and all $G$ and $i$ as in the statement of the proposition. Using the fibration sequences
$F_{d+1}\mathbb{H}^{alg}(X)\to F_{d}\mathbb{H}^{alg}(X)\to Gr_{d}\mathbb{H}^{alg}(X)\qquad J_{\geq n-d-1}(X)\to J_{\geq n-d}(X)\to J_{d}(X)[d]$
and the long exact sequences obtained after applying [,G[i]][-,G[i]], we see that by induction on $d$ we are thus reduced to show that, the induced morphism
$Ext^{i}_{naive}(J_{d}(X)[d],G)\to[Gr_{d}\mathbb{H}^{alg}(X),G[i]]$
is an isomorphism (for all $d$, and for all $G$ and $i$ as in the statement of the proposition). By the proposition 3.2, the right hand side can be identified with local cohomology $\bigoplus_{x\in X^{(n-j)}}H^{d+i}_{x}(X,G)$. We also note that by construction $J_{d}(X)=\prod_{x\in X^{(n-d)}}J_{x}(X)$, and thus we are reduced to show that for a given point $x\in X$, of codimension $d$, and all $i\geq 0$, the natural morphism
@@
is bijective under the extra condition stated in the proposition: $G$ is affine if $d+i>0$ and unipotent if $d+i>1$. Recall here that $J_{x}(X)$ corepresents the functor $G\mapsto H^{d}_{x}(X,G)$, where $d$ is the codimension of $x$ in $X$.
We first deal with the two special cases $i+d=0$ and $i+d=1$. When $i+d=0$, and thus $i=d=0$, we have $Ext^{0}(J_{x}(X),G)=Hom(J_{x}(X),G)\simeq H^{d}_{x}(X,G)$ and it is then obvious that $\psi^{0}_{G}$ is bijective. Suppose now that $d+i=1$, and thus $G$ is assumed to be affine. If $x$ is of codimension $1$, we have $i=0$, and again $\psi_{G}^{0}$ is bijective by definition of $J_{x}(X)$. If $x$ is of codimension 0 and $i=1$, we have to prove that the natural morphism
@@
where $K(X)$ is the fraction field of $X$. To show this we start by the following lemma treating the cases of $G=\mathbb{G}_{a}$ and $G=\mathbb{G}_{m}$.
@@
With the notation above the morphism
$\psi_{G}^{1}:Ext^{1}_{naive}(J_{0}(X),\mathbb{G}_{a})\to Ext^{1}_{naive}(J_{0}(X),\mathbb{G}_{m})\simeq 0$
is bijective.
Proof of the lemma. Suppose that we are given an extension of commutative pro-algebraic groups ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{G}_{a}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${J_{0}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$ By the universal property of $J_{0}(X)$, constructing a section $J_{0}(X)\to H$ is equivalent to give a point $s\in H(K(X))$, which projects to the universal element in $J_{0}(X)(K(X))$. But we have an exact sequence in fppffppf cohomology of $K(X)$ $H(K(X))\to J_{0}(X)(K(X))\to H^{1}_{fppf}(K(X),\mathbb{G}_{a})=0$, showing that $H(K(X))\to J_{0}(X)(K(X))$ is surjective. Therefore the short exact sequence splits showing that $Ext^{1}_{naive}(J_{0}(X),\mathbb{G}_{a})=0$. The proof for $\mathbb{G}_{m}$ is similar. $\Box$
@@
$0\to G\to\mathbb{G}_{m}\to\mathbb{G}_{m}\to 0\qquad 0\to G\to\mathbb{G}_{a}\to\mathbb{G}_{a}\to 0.$
We then consider the action of $\psi_{G}^{1}$ on the associated long exact sequences, and the lemma 3.8 shows that indeed $\psi_{G}^{1}$ is bijective for any $G$.
We now turn to the case where $i+d>1$ and $G$ is unipotent, which is proven in a very similar fashion than the case $i=1$ and $d=0$. To begin with, we prove that $\psi_{G}^{i}$ is an isomorphism for $G=\mathbb{G}_{a}$. Indeed, when $i=0$ this is obvious by the definition of $J_{x}(X)$. For $i>0$ the domain and codomain of $\psi_{\mathbb{G}_{a}}^{i}$ vanish. Indeed, $H^{d+i}_{x}(X,\mathbb{G}_{a})\simeq H^{d+i}_{x}(X,\mathcal{O})\simeq 0$ for $i>0$ by local duality for coherent sheaves. On the other hand, $Ext^{i}_{naive}(J_{x}(X),\mathbb{G}_{a})=0$ by the following lemma.